OpenStudy (anonymous):

A tennis ball machine serves a ball vertically into the air from a height of 2 feet, with an initial speed of 130 feet per second. Which equation correctly models the ball's height as a function of time?

OpenStudy (anonymous):

h(t)=-16^2 + 130t + 2 h(t)=-16^2 -130t + 2 h(t)=-16^2 + 2t + 130 h(t)=-16^2 - 2t +130

OpenStudy (anonymous):

h = -16t^2 + 130t + 2

OpenStudy (anonymous):

Thankss